FIND THE SUM OF ALL NUMBERS GREATER THAN 10,000 FORMED BY USING THE DIGITS 0,2,4,6,8 NO DIGIT BEING REPEATED IN ANY NUMBER.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear shaurya

Let 2 at first position

    rest 4 position can be filled by 4*4*4*4=256 ways

  every digit from 0,4,6,8 come at unit place 256/4 =64 times

 

 so some of all the digit at unit place = 64*(0+4+6+8)

each of 0,4,5,8 digit also occur at 64 times 10th hundred and thousand place

so sum of digit =64(0+4+6+8)(1+10+100+1000)   + 2 * 10000

similarly for the other digit

4 at first place sum =64(0+2+6+8)(1+10+100+1000)   + 4 * 10000

6 at first place sum =64(0+2+4+8)(1+10+100+1000)   + 6 * 10000

8 at first place sum =64(0+2+6+4)(1+10+100+1000)   + 8 * 10000

 

so total sum  =64(2+6+4+8)*3*(1+10+100+1000)   + (2+4+6+8) * 10000

             solve it this will be your final result


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin